OpenStudy (anonymous):

Complete the square y^2+1/3y=2/3

OpenStudy (anonymous):

half (1/3) and you get 1/6

OpenStudy (anonymous):

square (1/6) and you get 1/36

OpenStudy (anonymous):

add 1/36 to both sides of the equation

OpenStudy (anonymous):

y^2 +(1/3)y +(1/36) = 2/3 + 1/36 collect the LHS as a perfect sqaure

OpenStudy (anonymous):

(y+(1/6) ) ^2 = 25 / 36

OpenStudy (anonymous):

\[(y + \frac{1}{6})^2 = \frac{25}{36} \]